OpenStudy (anonymous):

Solve each system by elimination. I want to know how to do it, so show me each and every step so I will know how to do it for other problems ok? So thank you.

OpenStudy (anonymous):

4/3 = x - 8/9y 0=24x-51-15y

OpenStudy (anonymous):

mertsj

OpenStudy (anonymous):

Mertsj